find the area of the largest circle which can be cut from a square with an edge of 8cm.what is the area of the materials wasted?
Answer by Edwin McCravy(20086)

          

As you can see the green line segment is the diameter of the circle and it
is the same length as the edge of the square, so the diameter of the circle
is also 8 cm.  Since the radius of the circle is one-half of the diameter
the radius of the circle is 4cm.

Area of square = side² = (8cm)² = 64 cm²

Area of circle = pr² = (3.14)(4cm)² = (3.14)(16cm²) = 50.24 cm²

Area of waste =  Area of square - Area of circle = (64 cm²)-(50.24 cm²) = 13.76 cm²

Edwin